Grade 11IIT JEE Entrance Exam

Through the vertex O of a parabola y2=4x, chords OP & OQ are drawn at right angles to one another. Show that for all positions of P, PQ cuts the axis of the parabola at a fixed point. Find the fixed point and the locus of the middle point of PQ.

agamya goyal
10 Years agoGrade 11

1 Answer

Anoopam Mishra
10 Years ago
Let P be (t^2_{1},2t_{1}) and Q be (t^2_{2},2t_{2}).
The slope of line OP is 2/t_{1} and that of OQ is 2/t_{2}.
Since they are at right angle then 2/t_{1} * 2/t_{2} = -1 and hence t_{1} * t_{2} = -4.  
The coordinates of Q become (16/t^2_{1}, -8/t_{1}).
The slope of the line PQ is (2t_{1}+8/t_{1})/(t^2_{1}-16/t^2_{1}) = 2t_{1}/t^2_{1}-4.
The equation of line PQ is y-2t_{1} = 2t_{1}/t^2_{1}-4 \ (x-t^2_{1}).
Substituting y=0, we get x=4.
 
Now the second part of the question, let the midpoint of PQ be (h,k).
Then, 2h = t^2_{1}+16/t^2_{1} \ and \ 2k = 2t_{1}-8/t_{1}
4k^2 = 4t^2_{1}+64/t^2_{1}-32 = 4(2h)-32.
=> k^2 = 2(h-4)
Hence the locus of the midpoint of PQ is y^2 = 2(x-4).